Find two fraction between 3/7 and 3/4 and why it is those two fractions?
Answer by Alan3354(69443)

3/7 = 12/28
3/4 = 21/28
------------
13, 14, 15, 16.../28 are numbers between.
There is an infinite number between.
There's no "why"